HETAV PATEL

Grade 11,

Find the value of the polynomial 5x – 4x 2 + 3 at x = 2 and x = –1

 Find the value of the polynomial 5x – 4x2 + 3 at x = 2 and x = –1

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let the polynomial be f(x) =5x – 4x^2+ 3

Now, for x = 2,

f(2) = 5(2)– 4(2)^2+ 3

=> f(2) = 10 – 16 + 3 = –3

thus, value of the polynomial 5x – 4x^2+ 3 at x = 2 is -3.

Similarly, for x = –1,

f(–1) = 5(–1) – 4(–1)^2+ 3

=>f(–1) =–5–4 + 3 = -6

Thus, value of the polynomial at x = -1 is -6.


Thanks